# Exploring Citation Diversity in Scholarly Literature: An Entropy-Based Approach

### v1

@@ IV.1.1 Distribution pattern of citation counts of worldwide top institutes
Initially, we examine the distribution pattern of total citation counts $N_{cit}$ corresponding to a large number of institutes/universities ($N_{I}$) of various countries around the world. According to the Webomterics data (See Appendix for the data), considered for the analysis, rank 1 institute is the Harvard University of USA with citation count 27589889 and the Institute of Technology and Business of Czech Republic corresponds to rank 5661 with citation count 1004. The data furnishing a wide range of variation in $N_{cit}$. The distribution pattern (Fig. 1) provides valuable insights in understanding how the citation data is spread out or clustered around the world’s leading universities and institutes. Fig. 1(a) is the bar plot for the rank-size distribution, based on the ranks of the institutes $N_{I}$ and the corresponding size of citation counts $N_{cit}$; with an inset displaying the same plot in log-log scale for a clear understanding of the trend. Fig. 1(b) depicts the frequency distribution curve of total citation counts across different institutions in log-log scale. The distribution plot exhibits a power-law behavior in the higher citation end. The robustness of the fitted power law is checked by a goodness-of-fit test yielding satisfactory Kolmogorov-Smirnov distance $(KS)$ and $p$-value for the fit.
@@
Figure 1: Rank-size and frequency distribution plots of the total citation counts with the corresponding ranks of the institutes. The KS distance (KS), $p$-value and the exponent ($\alpha$) of the fitted power law distribution is given in the inset of (b).
To proceed further, we consider the citation data of top 10, 20, and 50 institutes or universities from each country across the globe. This data is found to be spread over 72, 55 and 25 countries respectively for top 10, 20 and 50 institutions. It is fascinating to note from Fig. 2 that the power law behaviour holds for all these 3 separate cases also. Each plot in the figure is accompanied with its respective KS distance $(KS)$ and $p$-value for the fit as well as the corresponding power law exponent $(\alpha)$. However, there is a variation in the value of the exponent with the change in the number of institutions considered for the analysis. The adherence of the consistent pattern of power law in the higher end of the citation counts indicates a predictable relationship between the rank of an institution and its citation count across different scales. Some recent studies have also demonstrated this power law trend in citation analyses [28, 29, 30].
@@
Figure 2: The citation distribution for top 10, 20 and 50 institutes in log-log scale for various countries across the world. Blue dots are the observed data points and red line represents the fitted power law with varying exponents ($\alpha$), KS distance (KS) and the $p$-values in each sub-plot.
@@ IV.1.2 Citation diversity in top institutes across the globe
This part of the discussion zooms on the diversity in the distribution of the citations of the top institutes or universities within each country. We employ the calculated D values, as previously explained in Section III, derived from the total citation count across all disciplines for each country’s top 10, 20, and 50 institutions. This approach allows us to classify the countries based on these diversity values (D) and also into 3 subgroups within each country based on high, medium and low citation counts ($N_{c}$). Tables 1,2 and 3 provide detailed breakdown of these grouping, respectively, for the top 10, 20 and 50 institutions across various nations; associated confidence intervals of the diversity measures are presented in Figs. 1,4 and 5, respectively, along with box-plot visualizations of the raw total citation data in each cases. Given the wide range of $N_{c}$ counts per country, we use a logarithmic scale for the y-axis in our box-plots (Figs. 3(b),4(b),5(b)) to effectively capture and represent the distribution of citation counts within different countries.
@@
It is evident from the $N_{c}$ count data of Table 1 that, within Group A, the USA stands out as the most highly cited country when examining its top 10 institutions. However, our analysis reveals a different leader in Group A, with Turkey emerging as the country with the highest citation diversity (lowest citation inequality). In Group B, while Switzerland emerges as the most frequently cited country, our study shows that Austria exhibits the highest citation diversity within the group. This indicates that although Switzerland may dominate in terms of citation volume, Austria’s citations are more evenly distributed. Conversely, Finland, despite being a part of the highly cited subgroup B1, registers the lowest citation diversity in Group B, suggesting a more concentrated citation pattern. Meanwhile, in Group D, Belgium stands out as the most frequently cited country, yet Norway surpasses it in citation diversity, indicating a wider spread of citations across top Norwegian institutes/universities.In all other groups and subgroups similar kind of results can be inferred. It is apparent that relying solely on total citation value or average citation counts fails to adequately appreciate the impression of citation analysis; the citation diversity measures are also required for a complete picture.
@@
By broadening our analysis to include the top 20 institutes, we have been able to study 55 countries across the globe as per the availability of data. In this case, we observe significant changes, compared to top 10 institutions, in the diversity measure (D) and the average citation count ($N_{c}$) for each country. The countries are again grouped as per their values of D and $N_{c}$ as in the case of top 10 institutes (Table 2 and Fig. 4). This expanded view makes the distinctions between countries based on these metrics more apparent. For instance, Israel is initially ranked in Group C with high D and maximum $N_{c}$ within this group when considering the top 10 institutes. However, when the scope is broadened to include the top 20 institutes, its performance metrics decline, moving it to Group E with a significantly lower D value. Similarly, the Netherlands is categorized in Group C with lower D and $N_{c}$ values when examining the top 20 institutes, but rises to Group A with much higher D and $N_{c}$ values when focusing on the top 10 institutes. In contrast, while considering the top 10 institutes, Morocco is positioned in Group A with a much higher D and $N_{c}$ values, but completely drops out of the rankings when the scope is expanded to the top 20 institutes.
@@
When we focus on the citation data for the worldwide top 50 institutes, we get data only on 25 countries whose diversity values are calculated from their total citations (Table 3 and Fig. 5). In the analysis, when focusing on the top 50 institutions, Taiwan and Thailand fall into Group E, characterized by lower D values. However, when considering only the top 20 institutions in these countries, they move to Group B, which has comparatively higher D values. Conversely, Spain is placed in Group A with high D and $N_{c}$ values when considering the top 20 institutions, but it shifts to Group B with lower D and $N_{c}$ values when the top 50 institutions are considered. In Group C, Canada is grouped with South Korea, Japan, and others, sharing similar D values but with significantly different $N_{c}$ values when considering the top 50 institutions. However, when focusing on the top 20 institutions, Canada moves to Group A, while South Korea and Japan are in Group B, with different D values.
@@
In conclusion, our novel citation diversity metric demonstrates its superiority over simple total citation counts in capturing a nation’s research landscape. We observed significant variations in diversity depending on different numbers of top institutes. For instance, Israel’s diversity dropped from $86.68\%$ (top 10) to $70.07\%$ (top 20), even falling out of the top 50 list altogether, indicating that the country has fewer than 50 renowned institutes. This highlights the crucial influence of a country’s concentration of high-performing institutes on its overall diversity score. Moreover, we observe that while the UK exhibits very high total citations across its top 10, 20, and 50 institutes, it is only classified in group A, characterized by a diversity range of 93% to 99%, when considering its top 10 and 20 institutes. However, when the top 50 institutes are taken into account, despite the high citation counts, the diversity value decreases, placing the UK in group B, with a diversity range of 89% to 93%. This indicates that although the UK maintains a strong citation performance, the citation diversity varies significantly with the number of institutes considered. When focusing on a smaller number of top institutes, the UK demonstrates a broader citation diversity, suggesting a wide-reaching influence of its most prominent research institutions. Conversely, India and USA displayed remarkable consistency in its diversity across all three institute tiers, suggesting a more balanced distribution of citations. However, expanding the scope to include more institutes reveals a drop in diversity, implying a more concentrated citation pattern. This highlights the importance of considering both citation count and diversity to fully understand the impact and reach of a country’s research output across different academic institutions across the globe. This analysis underscores the importance of considering citation distribution within a country, which total citation counts alone cannot effectively capture. Total citation counts often mask the underlying distribution of citations, potentially misleading interpretations. By employing our novel metric, we gain a clearer picture of how citations are distributed across a country’s top research institutions. Our approach provides a more nuanced understanding of a nation’s research landscape by revealing the distribution of citations amongst its leading institutions.
@@ IV.2 Citation diversity in various scientific disciplines
IV.2.1 Award winning scientists in recent times in different scientific disciplines
To develop a thorough understanding of citation diversity in different scientific disciplines, we implement our methodology from three distinct viewpoints. Table 4 showcases the calculated D values and $N_{c}$ counts per scientist for every discipline and Fig. 6 depicts these diversity measures, along with their 95% confidence intervals, and the distributions of individual numbers through their box-plots.
@@ IV.2.1 Award winning scientists in recent times in different scientific disciplines
It is noted that $N_{c}$ count per scientist in mathematics is minimum whereas its D value is not so low. On the other hand, the $N_{c}$ count per scientist in computer science is maximum but its D value is the lowest. So we can say that the citation inequality of award winners in computer science is very high as compared to the other disciplines. Additionally, we observe that the diversity of papers in computer science is relatively low compared to other subjects. However, the difference in paper diversity among various other subjects is not as pronounced as the difference in citation diversity between computer science and them.
@@ IV.2.2 Award winning scientists in old times in three principal disciplines
It is evident that, in earlier period, the diversity values for physics were the lowest in both the total citation and per-paper citation perspectives. However, in terms of the total publication perspective, physics had the highest diversity value. This suggests that citation inequality in physics was significant in previous times, whether considering the total citations or per-paper citations of award winners in this discipline. Conversely, the number of papers in physics was more evenly distributed compared to the other two subjects in historical time, as indicated by the higher diversity value for total publications. Additionally, the average citation count for chemistry was the lowest in both total citation and per-paper citation perspectives, while for the total publication, physics had the minimum average citation count. This again demonstrates that to obtain an exhaustive understanding of citation analysis, it is essential to look at the citation diversity values along with the citation counts of the publication data.
@@
From this comparison, it is evident that diversity values have increased across all disciplines from past eras to recent times, indicating a decrease in citation inequality. Notably, the recent era shows very small differences in diversity values among the three disciplines, highlighting a more even distribution of total number of citations, total number of papers, and per-paper citations among award winners across all disciplines. In contrast, these differences were much larger in the past. Additionally, we observe that while physics had the lowest diversity for the total citation in earlier times, it now has the highest D value compared to the other two disciplines in recent times. This suggests a significant improvement in the equality of citation distribution for physics. Again Fig. 8 reveals that the confidence interval for the total citation perspective is minimal, whereas the confidence intervals for the total publication and per-paper citation are comparatively large in both the recent and past eras. Thus, we can infer that the data for the total citation is more accurate compared to the other two perspectives. Overall, this comparison reveals that the citation distribution in physics has improved markedly in recent times compared to previous times.
IV.3 Citation diversity in the publication of Individual prize winning scientists across different disciplines
We now aspire to inspect the citation diversity in the publications of the individual prize winning scientists across various scientific disciplines. We have chosen the citation data for a total of 135 eminent scholars from the aforesaid scientific disciplines. We fixed on the Nobel prize winners in physics (30), chemistry (30), physiology/medicine (30) and economics (15), Abel prize winners in mathematics (15) and Turing award winners in computer science (15).
@@ IV.3.1 The Nobel Prize winners in Physics
The Nobel prize in physics has been awarded to 224 individuals between 1901 and 2023. For our investigation we have explored the citation counts of 30 Nobel laureates in physics, 15 from recent times (2019-2023) and 15 from the period (1901-1915). Using this data, we calculated the citation diversity values in the publication of these scientists following the methodology outlined above in Section III. Table 7 provides the citation diversity (D) values for each scientist considered along with their average citation counts ($N_{c}$). Fig. 9 illustrates the citation diversity values for each recent laureate, with Fig. 9(a) specifically highlighting the citation diversity of recent laureates. Notably, the confidence intervals for each point are very small, confirming the accuracy of these values. Additionally, Fig. 9(b) presents a box-plot of the citation counts of the laureates from raw citation data of their publications. The citation diversity values for the 15 recent laureates range from about 60% to 90%, with higher diversity correlating with lower average citation counts, underscoring the limitations of using average citations alone to represent a laureate’s citation distribution. Instead, the citation diversity measure offers a more precise depiction of these patterns. In Fig. 10, we observe the citation diversity values of earlier Nobel laureates, with Fig. 10(a) revealing a wide range of citation diversity from 20% to 80%. Larger confidence intervals further extend this range. Fig. 10(b) shows a box-plot based on their raw citation data, revealing that earlier laureates generally had lower citation counts but higher diversity values. This highlights how citation diversity accurately reflects citation distribution over time. Overall, the increase in diversity values from earlier to more recent laureates suggests a decline in citation inequality among Nobel laureates in physics over the years.

### v2

@@ IV.1.1 Distribution pattern of citation counts of worldwide top institutes
Initially, we examine the distribution pattern of total citation counts $N_{cit}$ corresponding to a large number of institutes/universities ($N_{I}$) from various countries around the world. According to the Webomterics data [1], considered for the analysis, rank 1 institute is the Harvard University of USA with citation count 27589889 and the Institute of Technology and Business of Czech Republic corresponds to rank 5661 with citation count 1004. The data furnishing a wide range of variation in $N_{cit}$. The distribution pattern (Fig. 1) provides valuable insights in understanding how the citation data is spread out or clustered around the world’s leading universities and institutes. Fig. 1(a) is the bar plot for the rank-size distribution, based on the ranks of the institutes $N_{I}$ and the corresponding size of citation counts $N_{cit}$; with an inset displaying the same plot in log-log scale for a clear understanding of the trend. Fig. 1(b) depicts the frequency distribution curve of total citation counts across different institutions in log-log scale. The distribution plot exhibits a power-law behavior in the higher citation end. The robustness of the fitted power law is checked by a goodness-of-fit test yielding satisfactory Kolmogorov-Smirnov distance $(KS)$ and $p$-value for the fit.
@@
Figure 1: Rank-size and frequency distribution plots of the total citation counts with the corresponding ranks of the institutes. The KS distance (KS), $p$-value and the exponent ($\nu$) of the fitted power law distribution is given in the inset of (b).
To proceed further, we consider the citation data of top 10, 20, and 50 institutes or universities from each country across the globe. This data is found to be spread over 72, 55 and 25 countries, respectively for top 10, 20 and 50 institutions. It is fascinating to note from Fig. 2 that the power law behaviour holds for all these 3 separate cases as well. Each plot in Fig. 2 is accompanied with its respective KS distance $(KS)$ and $p$-value for the fit as well as the corresponding power law exponent $(\nu)$. However, there is a variation in the value of the exponent with the change in the number of institutions considered for the analysis. The adherence of the consistent pattern of power law in the higher end of the citation counts [41] indicates a predictable relationship between the rank of an institution and its citation count across different scales. Some recent studies have also demonstrated this power law trend in citation analyses [4, 16, 23].
@@
Figure 2: The citation distribution for top 10, 20 and 50 institutes in log-log scale for various countries across the world. Blue dots are the observed data points and red line represents the fitted power law with varying exponents ($\nu$), KS distance (KS) and the $p$-values in each sub-plot.
@@ IV.1.2 Citation diversity in top institutes across the globe
Next we have studied the diversity in the distribution of the citations of the top institutes or universities within each country. We employ the calculated D values, as previously explained in Section III, derived from the total citation count across all disciplines for each country’s top 10, 20, and 50 institutions. This approach allows us to classify the countries based on these diversity values (D) and also into 3 subgroups within each group based on high, medium and low citation counts ($N_{c}$). Tables 1,2 and 3 provide detailed breakdown of these grouping, respectively, for the top 10, 20 and 50 institutions across various nations; associated confidence intervals of the diversity measures are presented in Figs. 1,4 and 5, respectively, along with box-plot visualizations of raw total citation data in each cases. Given the wide range of $N_{c}$ counts per country, we use a logarithmic scale for the y-axis in our box-plots (Figs. 3(b),4(b),5(b)) to effectively capture and represent the distribution of citation counts within different countries.
@@
It is evident from the $N_{c}$ count data of Table 1 that, within Group A, the USA stands out as the most highly cited country when examining its top 10 institutions. However, our analysis reveals a different leader in Group A, with Turkey emerging as the country with the highest citation diversity. In Group B, while Switzerland emerges as the most frequently cited country, our study shows that Austria exhibits the highest citation diversity within the group. This indicates that although Switzerland may dominate in terms of citation volume, Austria’s citations are more evenly distributed among its top 10 institutes. Conversely, Finland, despite being a part of the highly cited subgroup B1, registers the lowest citation diversity in Group B, suggesting a more concentrated citation pattern. Meanwhile, in Group D, Belgium stands out as the most frequently cited country, yet Norway surpasses it in citation diversity, indicating a wider spread of citations across top Norwegian institutes/universities. In all other groups and subgroups similar kind of results can be inferred. It is apparent that relying solely on total citation value or average citation counts fails to adequately appreciate the impression of citation analysis; the citation diversity measures are also required for a complete picture.
@@
By broadening our analysis to include the top 20 institutes, we have been able to study 55 countries across the globe as per the availability of data. In this case, we observe significant changes, compared to top 10 institutions, both in the diversity measure (D) and the average citation count ($N_{c}$) for each country. Countries are again grouped as per their values of D and $N_{c}$ as in the case of top 10 institutes (Table 2 and Fig. 4). This expanded view makes the distinctions between countries more apparent. For instance, Israel is initially ranked in Group C with high D and maximum $N_{c}$ within this group when considering the top 10 institutes. However, when the scope is broadened to include the top 20 institutes, its performance metrics decline, moving it to Group E with a significantly lower D value. Similarly, Netherlands is categorized in Group C with lower D and $N_{c}$ values when examining the top 20 institutes, but rises to Group A with much higher D and $N_{c}$ values when focusing on the top 10 institutes. These results imply that, in Finland and Netherlands, institutes ranked within 11 to 20 have significantly diverse and have lower citation counts compared to the top 10 institutes there, which were much more homogeneous in terms of citation counts. In contrast, while considering the top 10 institutes, Morocco is positioned in Group A with a much higher D and $N_{c}$ values, but completely drops out of the rankings when the scope is expanded to the top 20 institutes (as their is not many institutes outside the top 10 list in Morocco to have sizable/reportable citation data).
@@
When we focus on the citation data for top 50 institutes, we get data only on 25 countries whose diversity values are calculated from their total citations (Table 3 and Fig. 5). In the analysis, focusing on the top 50 institutions, Taiwan and Thailand fall into Group E, characterized by lower D values. However, when considering only the top 20 institutions in these countries, they move to Group B, which has comparatively higher D values. Conversely, Spain is placed in Group A with high D and $N_{c}$ values when considering the top 20 institutions, but it shifts to Group B with lower D and $N_{c}$ values when the top 50 institutions are considered. In Group C, Canada is grouped with South Korea, Japan, and others, sharing similar D values but with significantly different $N_{c}$ values when considering the top 50 institutions. However, when focusing on the top 20 institutions, Canada moves to Group A, while South Korea and Japan are in Group B, with different D values.
@@
In conclusion, our citation diversity metric complements total citation counts by providing additional insights that cannot be captured by citation counts alone. While total citations reflect overall research output and impact, the diversity metric highlights the evenness of the distribution of citations across different institutions, disciplines, and demographics. By considering both measures together, we gain a more comprehensive and nuanced understanding of a nation’s research landscape. In particular, this diversity index D ranges from 0 to 100 and can be interpreted as follows: a score of 100 would signify a perfectly even distribution of citations, while a score near 0 would indicate a highly uneven distribution. Therefore, for an example, Switzerland’s score of $D=89.18$ suggests that while the distribution is quite balanced, there is still some degree of unevenness which may be improved further (this unevenness is indeed more compared to the countries in Group A having values of $D>93$). The value of D are thus used for a basis of comparisons between countries in Tables 1,2 and 3, where countries in Group A have the most evenly distributed citations, while Group B includes countries with slightly less even citation distributions among their top institutes. The subsequent groups display progressively lower levels of evenness as the diversity scores decrease. We observed significant variations in diversity depending on different numbers of top institutes. For instance, Israel’s diversity decreased from $86.68\%$ (Group C) while considering top 10 institute to $70.07\%$ (Group E) while considering top 20 institute; the country is even dropped out of the top 50 list entirely, suggesting that the country has fewer than 50 renowned institutes, and the top 10 institutes are more homogeneous in terms of citation counts than the top 20 institutes there. This highlights the crucial influence of a country’s concentration of high-performing institutes on its overall diversity score. Moreover, we observe that while the UK exhibits very high total citations across its top 10, 20, and 50 institutes, it is only classified in group A, characterized by a diversity range of 93% to 99%, when considering its top 10 and 20 institutes. However, when the top 50 institutes are taken into account, despite the high citation counts, the diversity value decreases, placing the UK in group B, with a diversity range of 89% to 93%. This indicates that although the UK maintains a strong citation performance, the citation diversity varies significantly with the number of institutes considered. When focusing on a smaller number of top institutes, the UK demonstrates a broader citation diversity, suggesting a wide-reaching influence of its most prominent research institutions. Conversely, India and USA displayed remarkable consistency in its diversity across all three institute tiers, suggesting a more balanced distribution of citations. However, expanding the scope to include more institutes reveals a drop in diversity, implying a more concentrated citation pattern. This highlights the importance of considering both citation count and diversity to fully understand the impact and reach of a country’s research output across different academic institutions across the globe. Total citation counts often mask the underlying distribution of citations, potentially misleading interpretations. By employing our novel metric, we gain a clearer picture of how citations are distributed across a country’s top research institutions. Our approach provides a more nuanced understanding of a nation’s research landscape by revealing the distribution of citations amongst its leading institutions.
@@ IV.2 Citation diversity in various scientific disciplines
IV.2.1 Award winning scientists in recent times
To develop a thorough understanding of citation diversity in different scientific disciplines, we implement our methodology from three distinct viewpoints, namely total citations, total publications and per-paper citation count. Table 4 showcases the calculated D values and $N_{c}$ counts per scientist for every discipline and Fig. 6 depicts these diversity measures, along with their 95% confidence intervals, and the distributions of individual numbers through their box-plots.
@@ IV.2.1 Award winning scientists in recent times
It is noted that $N_{c}$ count per scientist in mathematics is minimum whereas its D value is not so low. On the other hand, the $N_{c}$ count per scientist in computer science is maximum but its D value is the lowest. So we can say that the total citation count of award winners in computer science is very high as compared to the other disciplines, but the diversity of papers in computer science is relatively low compared to other subjects. This indicated that, in computer science some award winners have excessively high number of papers and citations compared to some others. However, the number of papers and citation counts seems to be much more homogeneous across all winners in the other disciplines than in computer science. Additionally, the difference in paper and citation diversity among these other subjects is not as significant as the difference observed between computer science and them.
@@ IV.2.2 Award winning scientists in old times in three principal disciplines
It is evident that, in earlier period, the diversity values for physics were the lowest in both the total citation and per-paper citation perspectives. However, in terms of the total publication perspective, physics had the highest diversity value. This suggests that citation diversity in physics was significant in previous times, whether considering the total citations or per-paper citations of award winners in this discipline. Conversely, the number of papers in physics was more evenly distributed compared to the other two subjects in historical time, as indicated by the higher diversity value for total publications. Additionally, the average citation count for chemistry was the lowest in both total citation and per-paper citation perspectives, while for the total publication, physics had the minimum average citation count. This again demonstrates that to obtain an exhaustive understanding of citation analysis, it is essential to look at the citation diversity values along with the citation counts of the publication data.
@@
From this comparison, we can see that diversity values have increased across all disciplines from past eras to recent times. Notably, the recent era shows very small differences in diversity values among the three disciplines, highlighting a more even distribution of total number of citations, total number of papers, and per-paper citations among award winners across all disciplines. In contrast, these differences were much larger in the past. Additionally, we observe that while physics had the lowest diversity for the total citation in earlier times, it now has the highest D value compared to the other two disciplines in recent times, suggesting a significant improvement in the equality of citation distribution for physics. Again Fig. 8 reveals that the confidence interval for the total citation perspective is minimal, whereas the confidence intervals for the total publication and per-paper citation are comparatively large in both the recent and past eras. Thus, we can infer that the diversity estimate for the total citation is more reliable compared to the other two perspectives. Overall, this comparison reveals that the citation distribution in physics has improved markedly in recent times compared to previous times.
IV.3 Citation diversity in the publication of Individual prize winning scientists
We now aspire to inspect the citation diversity in the publications of the individual prize winning scientists across various scientific disciplines. We have chosen the citation data for a total of 135 eminent scholars from the aforesaid scientific disciplines. In particular, we have considered the Nobel prize winners in physics (30), chemistry (30), physiology/medicine (30) and economics (15), Abel prize winners in mathematics (15) and Turing award winners in computer science (15).
@@ IV.3.1 The Nobel Prize winners in Physics
The Nobel prize in physics has been awarded to 224 individuals between 1901 and 2023. For our investigation we have explored the citation counts of 30 Nobel laureates in physics, 15 from recent times (2019-2023) and 15 from the period (1901-1915). Using this data, we calculated the citation diversity values in the publication of these scientists following the methodology outlined above in Section III. Table 7 provides the citation diversity (D) values for each scientist considered along with their average citation counts ($N_{c}$). Fig. 9 illustrates the citation diversity values for each recent laureate, with Fig. 9(a) specifically highlighting the citation diversity of recent laureates. Notably, the confidence intervals for each point are very small, confirming the accuracy of these values. Additionally, Fig. 9(b) presents a box-plot of the citation counts of the laureates from raw citation data of their publications. The citation diversity values for the 15 recent laureates range from about 60% to 90%, with higher diversity correlating with lower average citation counts, underscoring the limitations of using average citations alone to represent a laureate’s citation distribution. In Fig. 10, we observe the citation diversity values of earlier Nobel laureates, with Fig. 10(a) revealing a wide range of citation diversity from 20% to 80%. Larger confidence intervals further extend this range. Fig. 10(b) shows a box-plot based on their raw citation data, revealing that earlier laureates generally had lower citation counts but higher diversity values. Overall, the increase in estimated diversity values from earlier to more recent laureates suggests a decline in citation diversity among Nobel laureates in physics over the years.